Plainwell Polymers is going to producing the main dashboard structure for a new car design. The dashboard will be produced in three colors with expected daily demand as follows:

Black: 545 units/day

Grey: 130 units/day

Tan: 210 units/day

The injection-molding machine can produce a total of 1,125 parts per day. What is the utilization of machine from these three parts as a percentage?
Answer #1

Given,

Production capacity of machine = 1125 parts/day

Total Daily demand = 545 + 130 + 210 = 885 parts/day

Hence, Utilization of machine = Total Daily demand/Production capacity of machine= 885 / 1125 = 78.67%.